Form a linear equation between
a) Age of your father and your age.
b) Age of your grandfather and father.
c) Your age and your mother's age

Answer by Theo(13342)

let x = your age
if your father is 25 years older than you, then x + 25 is your father's age.
if your mother is 20 years older than you, then x + 20 is your mother's age.
if your grandfather is 30 years older than your father, then x + 25 + 30 = x + 55 is your grandfather's age.
you have several equations that you can form.
y = x + 25 makes y equal to your father's age.
y = x + 20 makes y equal to your mother's age.
y = x + 55 makes y equal to your grandfather's age.
all of their ages are made in relationship to your age.
if you know your age, then you know their ages.
answers to your questions would be:

a) Age of your father and your age.
y = x + x + 25 which becomes y = 2x + 25.
if you are 20, then your father is 45 and the sum of your age and your father's age is 65.
y = 2x + 25 becomes y = 40 + 25 which becomes y = 65.

b) Age of your grandfather and father.
y = x + 25 + x + 55 which becomes y = 2x + 80.
if you are 20, then your father is 45 and your grandfather is 75 and the sum of their ages is 45 + 75 = 120.
y = 2x + 80 becomes y = 40 + 80 which becomes y = 120.

c) Your age and your mother's age
y = x + x + 20 which becomes y = 2x + 20.
if you are 20, then your mother is 40 and the sum of your age and her age is 60.
y = 2x + 20 becomes y = 40 + 20 which becomes y = 60.

the equation of y is solved in relationship to x, where x represents your age.
let me know if you have any questions.
theo